OpenStudy (anonymous):

Which fraction is equivalent to 72%? A.9/25 B.7/200 C.18/25 D.7 1/5

Mehek (mehek14):

72/100 simplify that 18/25

OpenStudy (anonymous):

numerator / denominator !! 9 / 25 = 0.36 7 / 200 = 0.035 18 / 25 = 0.72 1 / 5 = 0.2 just for reference, you can always do numerator / denominator to get a percentage, okay ??